# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ I.3.3 Resonant Inelastic x-ray Scattering
Resonant inelastic x-ray scattering (RIXS) is a photon scattering technique that has seen remarkable advancements over the past decade. RIXS is a two-step process involving photoelectric absorption and emission. As shown in Fig. 2(c), a core electron absorbs a photon and is excited into the valence band. This intermediate excited state then decays by emitting an x-ray photon. Due to its resonant nature, RIXS is described by the Kramers-Heisenberg cross-section Healy (1977); Ament et al. (2011), in contrast to the Thompson cross-section used for non-resonant IXS. Compared to IXS and INS, RIXS is sensitive to a wide range of quantum excitations, including spin, charge, orbital, lattice, and fractional quasiparticles. The intensity distribution in RIXS spectra also provides valuable insights into the strength of quasiparticle interactions. However, the RIXS cross-section is not directly proportional to the phonon dynamical structure factor. Additionally, despite significant progress, the state-of-the-art energy resolution of RIXS remains on the order of 20 meV, making it challenging to quantitatively analyze the phonon dynamical structure factor.
I.3.4 Momentum Resolved Electron Energy Loss Spectroscopy
@@
IXS and INS are generally bulk sensitive probes due to the deep penetration depth of hard x-rays and neutrons. For the interests in 2D materials and topological surface state, IXS and INS are limited by the small total scattering volume. To address this, high-resolution momentum resolved electron energy loss spectroscopy (HR-MEELS) has been developed as a powerful tool for investigating charge and phonon correlation functions on the surface Egerton (2011); Ibach and Mills (2013); Brydson (2020). Unlike the transmission EELS, MEELS experiments are typically conducted in reflection geometry with incident electron energy, $E_{i}$, less than 200 eV. By setting $E_{i}$=7.4 eV, energy and momentum resolutions that are better than 1 meV and 0.002Å-1 have been achieved, making MEELS a highly effective techniques for probing charge excitations in 2D materials. For electron scattering with small momentum transfer, $\bm{q}$, the HR-MEELS cross section can be formulated as:
$\frac{\mathrm{d}^{2}\sigma}{\mathrm{d}\Omega\mathrm{d}\omega}=\sigma_{0}V_{\textrm{eff}}^{2}(\bm{q})\int_{-\infty}^{0}z_{1}z_{2}e^{-|\bm{q}||dz_{1}+dz_{2}|}\cdot S(\bm{q},z_{1},z_{2},\omega),$ (26)
@@
I.3.5 Other Techniques
Numerous other techniques are also sensitive to phonon excitations. For instance, resonant ultrasound spectroscopy probes the quasi-elastic phonon excitations in the MHz range Migliori (2016). Additionally, pump-probe and time-domain spectroscopy Fushitani (2008); Fischer et al. (2016); Hangyo et al. (2005); Koch et al. (2023), utilizing tabletop laser setup or advanced free-electron lasers, provide access to coherent optical and acoustic phonons, as well as their interactions with spin and charge excitations. These methods offer versatile tools for exploring the dynamic behavior of phonons and their coupling to other quantum excitations in materials.
@@ II Topological Phonons: General Topological Band Theory
Topological band theory (TBT) is an extension of traditional band theory that incorporates the principles of topology. It has been widely applied to various condensed matter platforms, including electrons Fu et al. (2007); Fu and Kane (2007); Hasan and Kane (2010); Qi and Zhang (2011); Tokura et al. (2019); Murakami (2007); Burkov (2016); Zhang et al. (2019a); Vergniory et al. (2019b); Tang et al. (2019a); Lv et al. (2021); Bradlyn et al. (2017); Po et al. (2017), phonons Sheng et al. (2006); Zhang et al. (2010); Liu et al. (2017b, a, 2018); Zhang et al. (2018a, 2019b); Li et al. (2021b); Xu et al. (2024b), magnons Zhang et al. (2013); Owerre (2016); McClarty (2022), photons Lu et al. (2014, 2016a, 2016b); Ozawa et al. (2019), and artificial systems Kane and Lubensky (2014); Wang et al. (2015a); Süsstrunk and Huber (2015); Yang et al. (2015); Ma et al. (2019b); Zhang et al. (2019d); Huber (2016a); Nash et al. (2015); Khanikaev and Shvets (2017); Imhof et al. (2018); Luo et al. (2018); Ningyuan et al. (2015); Lu et al. (2015). In this section, we will first introduce the basics of topological classifications of gapped systems, which is not limited to electronic systems with an energy gap near the Fermi energy. Then, we will expand the topological band theory to phonon systems, which is the mainly focus in this review. We emphasize the different topological classifications for gapless and gapped phonon excitations and discuss their corresponding topological invariants under various symmetries. Since phonon excitations generally preserve time-reversal symmetry ($\mathcal{T}$), our discussion will focus primarily on time-reversal-invariant systems.
II.1 Topological Classifications in Gapped Phonon Bands
@@
The “ten-fold way” refers to a classification framework in condensed matter physics and quantum field theory, primarily used to categorize symmetry classes of Hamiltonians, particularly in the study of of topological insulators Fu and Kane (2007); Qi and Zhang (2011); Hasan and Kane (2010); Hasan and Moore (2011); Shen (2012); Bernevig (2013); Khanikaev et al. (2013); Ando (2013); Tokura et al. (2019) and topological superconductors Leijnse and Flensberg (2012); Ando and Fu (2015); Sato and Ando (2017). Such classification is grounded in the presence or absence of three foundational discrete symmetries, i.e., time-reversal symmetry ($\mathcal{T}$), particle-hole symmetry ($\mathcal{C}$), and chiral (or sublattice) symmetry ($\mathcal{S}$). The idea of the ten-fold way dates back at least to 1996, when Altland and Zirnbauer (1997) discovered that substances can be divided into 10 kinds Kitaev (2009); Ryu et al. (2010). Equations ((28)-(30)) provide the mathematical expressions for these three symmetry operators, as well as the transformation properties of the Hamiltonian under these symmetries.
$\mathcal{T}^{-1}H(k)\mathcal{T}$ $=H(-k),$ $\mathcal{T}$ $=U_{T}\mathcal{K},$ $U_{T}U_{T}^{*}$ =±$=\pm 1,$ (28)
$\mathcal{C}^{-1}H(k)\mathcal{C}$ $=-H(-k),$ $\mathcal{C}$ $=U_{C}\mathcal{K},$ $U_{C}U_{C}^{*}$ =±$=\pm 1,$ (29)
@@
In a system exhibiting time-reversal symmetry ($\mathcal{T}$), $\mathcal{T}^{2}$ = $-1$ or +1, corresponding to half-integer spin systems (including spinful electronic systems) and integer-spin systems (e.g phonons and photons), respectively. When the time-reversal symmetry is broken, phenomena like the quantum anomalous Hall effect (QAHE) emerge Haldane (1988); Liu et al. (2008); Chang et al. (2013); Liu et al. (2016); Deng et al. (2020); Chang et al. (2023), and this case is symbolically expressed as $\mathcal{T}^{2}$ = 0 (where “0” signifies the absence of time-reversal symmetry). Similarly, particle-hole symmetry ($\mathcal{C}$) is characterized by an anti-unitary operator with $\mathcal{C}^{2}$ = ±$\pm 1$, or $\mathcal{C}^{2}$ = 0 in the absence of particle-hole symmetry, yielding three possible cases. Chiral symmetry (or sublattice symmetry in certain contexts) can be interpreted as a combination of $\mathcal{T}$ and $\mathcal{C}$ ($\mathcal{S}$ = $\mathcal{TC}$), which also results in three possibilities, with $\mathcal{S}^{2}=\pm 1,0$. Thus, we have identified 9 distinct types of systems by considering the possible values of $\mathcal{T}^{2},\ \mathcal{C}^{2},\ \mathcal{S}^{2}$, each of which can take values of 1, $,\ 1$ 1 or 0. This naturally raises the question: what constitutes the tenth kind in the “ten-fold way”? The answer lies in systems that lack both $\mathcal{T}$ and $\mathcal{C}$ individually, yet still exhibit symmetry under their combined operation, $\mathcal{S}$ = $\mathcal{TC}$. In such systems, invariance is preserved when particles and holes are interchanged and time is reversed, even though neither symmetry exists alone. This unique scenario completes the ten-fold classification.
@@
Figure 3: “Cartan label” for “ten-fold way” classes of single-particle Hamiltonians $H$, based on time-reversal symmetry ($\mathcal{T}$), particle-hole symmetry ($\mathcal{C}$) and chiral symmetry ($\mathcal{S}$ = $\mathcal{TC}$). Adapted from Ref. Ryu et al. (2010).
@@
Figure 4: “Ten-fold way” table for topological insulators and superconductors as a function of spatial dimension $d$ and three discrete symmetries, such that the symmetry classes are arranged to reveal a periodic pattern in spatial dimensionality. The “Cartan label” (first column) is shown in Fig. 3. These ten classes are subsequently categorized into two groups: the complex and real ones, mainly contingent upon the complexity of the Hamiltonian. $\mathbb{Z}$ and $\mathbb{Z}_{2}$ represent the distinct phases within a specific symmetry class of topological insulators/superconductors, distinguished by an integer invariant. Within a specific $\mathbb{Z}$ group, the topological invariant is isomorphic to the $\mathbb{Z}$-type Abelian group and can assume any integer value. For the $\mathbb{Z}_{2}$ class, the topological invariant is isomorphic to the $\mathbb{Z}_{2}$-type Abelian group, such that the topological invariant can take two classes of values, the odd ones and the even ones. “0” means that no topological insulator/superconductor is present, indicating that all quantum ground states are topologically trivial. Adapted from Ref. Ryu et al. (2010)

### v2

@@ I.3.4 Resonant Inelastic x-ray Scattering
Resonant inelastic x-ray scattering (RIXS) is a photon scattering technique that has seen remarkable advancements over the past decade ament2011resonant ; RIXSReview . RIXS is a two-step process involving photoelectric absorption and emission. As shown in Fig. 2(c), a core electron absorbs a photon and is excited into the valence band. This intermediate excited state then decays by emitting an x-ray photon. Due to its resonant nature, RIXS is described by the Kramers-Heisenberg equations healy1977generalization ; ament2011resonant . The RIXS scattering matrix element is given by:
$M_{if}=\sum_{n}\frac{\langle f|\mathcal{D}_{\mathbf{k}^{\prime},\hat{\varepsilon}_{f}}^{\dagger}|n\rangle\langle n|\mathcal{D}_{\mathbf{k},\hat{\varepsilon}_{i}}|i\rangle}{E_{n}-E_{i}-\hbar\omega_{\mathbf{k}}+\mathrm{i}\Gamma_{n}/2}$ (30)
where $\Gamma_{n}/2$ is the inverse core-hole lifetime in units of energy, $E_{n}$ is the energy of the intermediate state. $\mathcal{D}_{\mathbf{k}^{\prime},\hat{\varepsilon}_{f}}^{\dagger}$ and $\mathcal{D}_{\mathbf{k},\hat{\varepsilon}_{i}}$ are the emission and absorption operators that are given by the dipole term in Eq. 22.
Compared to IXS and INS, RIXS is sensitive to a wide range of quantum excitations, including spin, charge, orbital, lattice, and fractional quasiparticles ament2011resonant ; RIXSReview . The intensity distribution in RIXS spectra also provides valuable insights into the strength of quasiparticle interactions. On the other hand, RIXS cross-section is not simply proportional to the phonon dynamical structure factor. Additionally, despite significant progress, the state-of-the-art energy resolution of RIXS remains on the order of 20 meV, making it challenging to quantitatively analyze the phonon dynamical structure factor.
I.3.5 Momentum Resolved Electron Energy Loss Spectroscopy
@@
IXS and INS are generally bulk sensitive probes due to the deep penetration depth of hard x-rays and neutrons. For the interests in 2D materials and topological surface state, IXS and INS are limited by the small total scattering volume. To address this, high-resolution momentum resolved electron energy loss spectroscopy (HR-MEELS) has been developed as a powerful tool for investigating charge and phonon correlation functions on the surface egerton2011electron ; ibach2013electron ; Vig2017Measurement ; brydson2020electron . Unlike the transmission EELS, HR-MEELS experiments are typically conducted in reflection geometry with incident electron energy, $E_{i}$, less than 200 eV. It has been shown that by setting $E_{i}<$=10 eV, HR-MEELS can achieve energy and momentum resolutions better than 1 meV and 0.002Å-1, making HR-MEELS a highly effective techniques for probing phonon and charge excitations in 2D materials. For electron scattering with small momentum transfer, $\bm{q}$, the HR-MEELS cross section can be formulated as:
$\frac{\mathrm{d}^{2}\sigma}{\mathrm{d}\Omega\mathrm{d}\omega}=\sigma_{0}V_{\textrm{eff}}^{2}(\bm{q})\int_{-\infty}^{0}dz_{1}dz_{2}e^{-|\bm{q}||z_{1}+z_{2}|}\cdot S(\bm{q},z_{1},z_{2},\omega),$ (31)
@@
I.3.6 Other resonant and stimulating Techniques
Phonons can also be detected by other resonant and stimulating techniques. For instance, resonant ultrasound spectroscopy probes quasi-elastic phonon excitations in the MHz range migliori2016resonant . This technique complements experimental probes described in previous sections that typically focus on phonons in THz-range. Additionally, pump-probe and time-domain spectroscopy fushitani2008applications ; fischer2016invited ; hangyo2005terahertz ; koch2023terahertz , utilizing tabletop laser setup or advanced free-electron lasers, provide access to coherent optical and acoustic phonons, as well as their interactions with spin and charge excitations. These methods offer versatile tools for exploring the dynamic behavior of phonons and their coupling to other quantum excitations in materials.
@@ II Topological Phonons: General Topological Band Theory
Topological band theory (TBT) is an extension of traditional band theory that incorporates the principles of topology. It has been widely applied to various condensed matter platforms, including electrons fu2007topological ; hasan2010colloquium ; qi2011topological ; tokura2019magnetic ; murakami2007phase ; burkov2016topological ; zhang2019catalogue ; vergniory2019complete ; tang2019comprehensive ; lv2021experimental ; bradlyn2017topological ; po2017symmetry , phonons sheng2006theory ; Zhang2010Nov ; liu2017model ; liu2017pseudospins ; liu2018berry ; zhang2018double ; zhang2019phononic ; li2021computation ; xu2024catalog , magnons zhang2013topological ; owerre2016first ; mcclarty2022topological , photons lu2014topological ; lu2016symmetry ; lu2016topological ; ozawa2019topological , and artificial systems kane2014topological ; wang2015topological ; susstrunk2015observation ; yang2015topological ; ma2019topological ; zhang2019second ; huber2016topological ; nash2015topological ; khanikaev2017two ; imhof2018topolectrical ; luo2018topological ; ningyuan2015time ; lu2015experimental . In this section, we will start by introducing the basics of topological classifications of gapped systems. We will then expand the topological band theory to phonon excitations that is the mainly focus of this review. We emphasize the different topological classifications for gapless and gapped phonon excitations and discuss their corresponding topological invariants under various symmetries. Since phonon excitations generally preserve time-reversal symmetry ($\mathcal{T}$), our discussion will focus primarily on time-reversal-invariant systems.
II.1 Comparison Between Different Topological Quasi-particles
Topological phonons, grounded in the principles of band theory, exhibit parallels with other topological quasiparticles, particularly those within spinless electronic band structures. They share characteristics such as various types of Weyl points, Dirac points, and nodal lines/rings, along with the emergence of topological surface states due to the “bulk-surface correspondence”. In electronic systems without spin-orbit coupling, the spin degree of freedom is decoupled from other degrees of freedom. Therefore, the spin system has the full SU(2) symmetry. After decoupling the spin degree of freedom, we apply the topological band theory for spinless electronic systems. Phonons, being spin-0 bosons, can be treated within the same topological framework as spinless electronic systems. Topological states of phonons can be defined both in the presence and absence of time-reversal symmetry ($\mathcal{T}$). However, we note that breaking $\mathcal{T}$-symmetry in phonon spectra is uncommon and typically requires specific conditions—such as external magnetic fields Baydin2022_PRB ; Cheng2020_Cd3As2 ; Luo2023_CeF3_Science ; Chaudhary2024_PRB ; Wu2023_Fe2Mo3O8_NP ; David2024_CoTiO3_PNAS ; Nova2017_ErFeO3_NP ; Schaack1977_CeCl3 ; CeCl3_2022 ; Niu2021_phononMag ; Xue2025_Extrinsic , coupling to magnons Ren2024_PH_SPIN_PRX ; Liu2021_Magnon_phonon , or intrinsic magnetic order CoSnS_CP_2025 ; CoSnS_PRL_2025 ; zhang2025electronic . Nonetheless, the fundamental distinctions between phonons and electrons could pave the way for the revelation of novel topological phenomena unique to phonons. Thus, topological phonons are endowed with distinctive attributes that set them apart from their counterparts among other topological quasiparticles.
One difference between fermionic electrons and bosonic phonons lies in the robustness of gapless topological states such as nodal-line or nodal-ring states. In electronic systems, these states are often fragile and can be gaped by spin-orbit coupling (SOC), leading to either trivial or topological insulating states. This is, however, not the case for phonons as phonons are spin-0 quasiparticles with orbital degree of freedom, hence precludes SOC effect zhang2019phononic . Furthermore, in three dimensions, the threefold degeneracy of acoustic phonons at the $\Gamma$ point, arising from their nature as Goldstone modes, becomes spin-1 Weyl phonons in chiral crystals, where the degeneracy between the two transverse modes is lifted. Phonons can possess additional quantum degrees of freedom, such as pseudospin and angular momentum, which endow them with novel physical properties associated with their topological states zhang2023weyl .
II.2 Topological Classifications in Gapped Phonon Bands
@@
The “ten-fold way” refers to a universal classification framework in condensed matter physics and quantum field theory, primarily used to categorize the symmetry classes of Hamiltonians, particularly in studies of topological insulators fu2007topological ; qi2011topological ; hasan2010colloquium ; hasan2011three ; shen2012topological ; bernevig2013topological ; khanikaev2013photonic ; ando2013topological ; tokura2019magnetic and topological superconductors leijnse2012introduction ; ando2015topological ; sato2017topological . This classification is determined by the presence or absence of three fundamental discrete symmetries, i.e., time-reversal ($\mathcal{T}$), particle-hole ($\mathcal{C}$), and chiral (or sublattice, $\mathcal{S}$) symmetry, as expressed in Eqs. (33)–(35). The concept of the ten-fold way dates back to the work of Altland and Zirnbauer Altland1997 , who showed that physical systems can be categorized into ten distinct symmetry classes kitaev2009periodic ; ryu2010topological .
Mathematically, the ten-fold classification is grounded in K-theory, which provides a systematic framework for classifying gapped Hamiltonians under given symmetry constraints. Within this framework, the symmetry class determines the associated “classifying space”, while the Clifford algebra leads to a repeating pattern in the classification, also known as “Bott periodicity”. Specifically, an eightfold periodicity for real classes and a twofold periodicity for complex ones. This periodic structure results from the recursive representation properties of Clifford algebra, yielding periodic recurrences of topological invariants such as $\mathbb{Z}$, $\mathbb{Z}_{2}$, or 0, depending on spatial dimension.
Physically, the ten-fold classification establishes a direct correspondence between the symmetry and dimensionality of the system. This framework also uncovers the existence of quantized bulk invariants and symmetry-protected gapless boundary states. Therefore, the ten-fold classification table provides a unified theoretical foundation that connects diverse topological phenomena, ranging from the integer quantum Hall effect to topological superconductivity.
Consider a Hamiltonian, $\mathcal{H}(k)$, whose transformations under the symmetry operations $\mathcal{T}$, $\mathcal{C}$, and $\mathcal{S}$ are given by:
$\mathcal{T}^{-1}H(k)\mathcal{T}$ $=H(-k),$ $\mathcal{T}$ $=U_{T}\mathcal{K},$ $U_{T}U_{T}^{*}$ $=\pm 1,$ (33)
$\mathcal{C}^{-1}H(k)\mathcal{C}$ $=-H(-k),$ $\mathcal{C}$ $=U_{C}\mathcal{K},$ $U_{C}U_{C}^{*}$ $=\pm 1,$ (34)
@@
In a system exhibiting $\mathcal{T}$, $\mathcal{T}^{2}$ = 1-1 or +1 are respectively corresponding to half-integer spin systems (including spinful electronic systems) and integer-spin systems (e.g phonons and photons), respectively. When the time-reversal symmetry is broken, phenomena like the quantum anomalous Hall effect (QAHE) emerge haldane1988model ; liu2008quantum ; chang2013experimental ; liu2016quantum ; deng2020quantum ; chang2023colloquium , and this case is symbolically expressed as $\mathcal{T}^{2}$ = 0 (where “0” signifies the absence of time-reversal symmetry). Similarly, particle-hole symmetry ($\mathcal{C}$) is characterized by an anti-unitary operator with $\mathcal{C}^{2}$ = $\pm 1$, or $\mathcal{C}^{2}$ = 0 in the absence of particle-hole symmetry, yielding three possible cases. Chiral symmetry (or sublattice symmetry in certain contexts) can be interpreted as a combination of $\mathcal{T}$ and $\mathcal{C}$ ($\mathcal{S}$ = $\mathcal{TC}$), which also results in three possibilities, with $\mathcal{S}^{2}=\pm 1,0$. Thus, we have identified 9 distinct types of systems by considering the possible values of $\mathcal{T}^{2},\ \mathcal{C}^{2},\ \mathcal{S}^{2}$, each of which can take values of 1, 1-1,\ 1 or 0. This naturally raises the question: what constitutes the tenth kind in the “ten-fold way”? The answer lies in systems that lack both $\mathcal{T}$ and $\mathcal{C}$ individually, yet still exhibit symmetry under their combined operation, $\mathcal{S}$ = $\mathcal{TC}$. In such systems, invariance is preserved when particles and holes are interchanged and time is reversed, even though neither symmetry exists alone. This unique scenario completes the ten-fold classification.
@@
Table 1: “Cartan label” for “ten-fold way” classes of single-particle Hamiltonians $H$, based on time-reversal symmetry ($\mathcal{T}$), particle-hole symmetry ($\mathcal{C}$) and chiral symmetry ($\mathcal{S}$ = $\mathcal{TC}$). Adapted from Ref. ryu2010topological .
[…]
A (unitary) 0 0 0 $U(N)$ $U(2n)/U(n)\times U(n)$
AI (orthogonal) +1 0 0 $U(N)/O(N)$ $Sp(2n)/Sp(n)\times Sp(n)$
AII (symplectic) -1 0 0 $U(2N)/Sp(2N)$ $O(2n)/O(n)\times O(n)$
AIII (ch. unit.) 0 0 1 $U(N+M)/U(N)\times U(M)$ $U(n)$
BDI (ch. orth.) +1 +1 1 $O(N+M)/O(N)\times O(M)$ $U(2n)/Sp(2n)$
CII (ch. sympl.) -1 -1 1 $Sp(N+M)/Sp(N)\times Sp(M)$ $U(2n)/O(2n)$
D (BdG) 0 +1 0 $SO(2N)$ $O(2n)/U(n)$
C (BdG) 0 -1 0 $Sp(2N)$ $Sp(2n)/U(n)$
DIII (BdG) -1 +1 1 $SO(2N)/U(N)$ $O(2n)$
CI (BdG) +1 -1 1 $Sp(2N)/U(N)$ $Sp(2n)$
@@
Table 2: “Ten-fold way” table for topological insulators and superconductors as a function of spatial dimension $d$ and three discrete symmetries, such that the symmetry classes are arranged to reveal a periodic pattern in spatial dimensionality. The “Cartan label” (first column) is shown in Fig. 1. These ten classes are subsequently categorized into two groups: the complex and real ones, mainly contingent upon the complexity of the Hamiltonian. $\mathbb{Z}$ and $\mathbb{Z}_{2}$ represent the distinct phases within a specific symmetry class of topological insulators/superconductors, distinguished by an integer invariant. Within a specific $\mathbb{Z}$ group, the topological invariant is isomorphic to the $\mathbb{Z}$-type Abelian group and can assume any integer value. For the $\mathbb{Z}_{2}$ class, the topological invariant is isomorphic to the $\mathbb{Z}_{2}$-type Abelian group, such that the topological invariant can take two classes of values, the odd ones and the even ones. “0” means that no topological insulator/superconductor is present, indicating that all quantum ground states are topologically trivial. Adapted from Ref. ryu2010topological
$d$
Cartan 0 11 22 33 44 55 66 77 88 99 $10$ $11$ $\dots$
Complex case:
A $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\dots$
AIII 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ $\dots$
Real case:
AI $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 $\dots$
BDI $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0
D $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 $\dots$
DIII 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ $\dots$
AII $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\dots$
CII 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\dots$
C 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}$ 0 $\dots$
CI 0 0 0 $2\mathbb{Z}$ 0 $\mathbb{Z}_{2}$ $\mathbb{Z}_{2}$ $\mathbb{Z}$ 0 0 0 $2\mathbb{Z}$ $\dots$